A 200,000 kg train moving at 3m/s hits a initally still train that is 600,000 kg , they link t ogeatjer and keep moving, how fast are they moving afterward
Answer #1

By conservation of momentum we have

m1u1+m2u2=m1v1+m2v2

20020*3=(600000+20020)v

v=0.096 m/sec

they moving afterward with 0.096 m/sec
Answer #3

no external forces act

so,momentum will conserv

initial momentum=final momentum

initiam momentum=200000*3=600000kg.m/s

final momentum=(200000+600000)*V

initial momentum=final momentum

600000=800000*V

V=6/8=3/4=0.75m/sec